OpenStudy (anonymous):

Lauren has gross pay of $765 and federal tax withholdings of $68. Determine Lauren's net pay if she has the additional items withheld: Social Security tax that is 6.2% of her gross pay Medicare tax that is 1.45% of her gross pay State tax that is 22% of her federal tax a. $691.52 b. $666.21 c. $636.98 d. $623.52

OpenStudy (anonymous):

screw the government and their tax law!

OpenStudy (triciaal):

gross =768 fed = 68 ss = 6.2/100 times 765 med = 1.45/100 times 765 state = 22/100 times 68 gross - ( sum of all deductions) = net pay

OpenStudy (mathstudent55):

Start with $765 Subtract $68 Subtract 6.2% of $765 = 0.062 * $765 = <--- fill this in Subtract 1.45% of $765 = 0.0145 * $765 = <--- fill this in Subtract 22% of $68 = 0.222 * $65 = <--- fill this in -------------------------------------------- Net pay <--- subtract all those amounts from $765 and write the result here